### v1

# Mapping the Trust Terrain: LLMs in Software Engineering - Insights and Perspectives

Abstract.
Applications of Large Language Models (LLMs) are rapidly growing in industry and academia for various software engineering (SE) tasks. As these models become more integral to critical processes, ensuring their reliability and trustworthiness becomes essential. Consequently, the concept of trust in these systems is becoming increasingly critical. Well-calibrated trust is important, as excessive trust can lead to security vulnerabilities, and risks, while insufficient trust can hinder innovation. However, the landscape of trust-related concepts in LLMs in SE is relatively unclear, with concepts such as trust, distrust, and trustworthiness lacking clear conceptualizations in the SE community. To bring clarity to the current research status and identify opportunities for future work, we conducted a comprehensive review of $88$ papers: a systematic literature review of $18$ papers focused on LLMs in SE, complemented by an analysis of $70$ papers from broader trust literature. Additionally, we conducted a survey study with $25$ domain experts to gain insights into practitioners’ understanding of trust and identify gaps between existing literature and developers’ perceptions. The result of our analysis serves as a roadmap that covers trust-related concepts in LLMs in SE and highlights areas for future exploration.
Trust, Distrust, Trustworthiness, LLMs
1. Introduction
The use of LLMs in SE is becoming increasingly prevalent, with applications spanning bug fixing  (Tufano et al., 2018), defect prediction  (Wang et al., 2016), input generation  (Liu et al., 2017), and various other tasks  (Gu et al., 2018a; Guo et al., 2017; Corley et al., 2015; Liu et al., 2018; Watson et al., 2021; Gu et al., 2018b). These models are now widely adopted by software developers, researchers, and students, demonstrating their potential to enhance productivity and automate complex SE tasks.
However, the successful integration of LLMs into SE workflows depends not only on their technical capabilities but also on how practitioners perceive and trust them (Roychoudhury et al., 2025). For example, misaligned trust between LLM and practitioner, either excessive (i.e., over-trust) or insufficient (i.e., under-trust), can significantly impact the effectiveness and security of these LLMs. Achieving the right alignment is crucial, as both extremes entail undesirable consequences: over-trust in LLMs may lead to security vulnerabilities, data integrity risks, and erroneous decision-making  (Thorne, 2024), while under-trust can hinder adoption, reducing the potential benefits of LLMs  (Balayn et al., 2024).
The alignment of trust between LLMs and practitioners determines the myriad of feasible trust relations within software engineering settings. Trust relations are characterized not only by the trustor, defined as the party that decides whether or not to trust, and the trustee, defined as the party or the object to be trusted  (Harjit Sekhon and Devlin, 2014), but also by a specific activity  (Balayn et al., 2024). For example, a typical trust relation scenario is when a practitioner (i.e., trustor) accepts or rejects a snippet generated by an LLM such as GitHub Copilot (i.e., trustee). Furthermore, the perception of trust varies among the trustor, SE-tasks, and domains (e.g., SE, HCI, DL) within these relationships (Balayn et al., 2024). For example, a developer may confidently rely on an LLM to generate boilerplate code but hesitate to use it for test case generation due to concerns about domain-specific accuracy. Understanding the core concepts of trust in these contexts is crucial for uncovering the complexities of these relationships.
However, core concepts such as trust, distrust, and trustworthiness are often used interchangeably, obscuring their precise meanings  (Sharma et al., 2024). As a result, fundamental questions about trust in LLM4SE remain unresolved: What factors shape trust, and do these factors vary across experience levels, such as between expert and novice developers, or across different SE tasks? Is distrust simply the absence of trust, or does it represent a distinct state? Addressing these unresolved questions is crucial for adopting LLMs in SE workflows successfully, as demonstrated by research findings in the field of Human-Computer Interaction (HCI) (Choung et al., 2022; Glikson and Woolley, 2020). These HCI studies have shown that trust is a crucial factor that influences the willingness of stakeholders to adopt LLMs. Unfortunately, existing studies provide a limited foundational understanding of how trust is defined, the factors that influence it, and the metrics used to assess it.
To establish this foundational understanding, we must first assess the current state of research on trust in LLM4SE. Nonetheless, a comprehensive review of existing work has not been conducted, making it difficult to determine what aspects of trust have been explored and where significant gaps remain. The absence of a structured overview of trust concepts hinders efforts to define key trust-related concepts, identify influencing factors, and develop reliable evaluation methodologies. Additionally, the SE practitioners’ perspective on the definition, factors, and metrics of trust remains largely under-explored, leaving a disconnect between academic research and real-world needs. Without a clear understanding of the existing landscape, advancing trust-aware LLM systems for SE remains challenging.
To address these challenges, we conducted a comprehensive approach that combines literature review with practitioner insights. First, we reviewed $18$ SE-focused papers to examine how trust-related concepts are defined, what factors influence trust in LLM4SE, and what metrics are used for evaluation. Second, we conducted a survey study with $25$ SE practitioners to capture real-world perspectives on trust in LLM4SE, identify task-specific trust factors, and explore variations in trust perceptions at different levels of expertise. In addition, we carried out a complementary analysis of $70$ papers from disciplines such as DL, HCI, and automation to enrich and complement the trust conceptualization from other fields.
By integrating findings from the software engineering literature and practitioner experiences, we offer a comprehensive perspective on the current state of trust research in LLM4SE. Beyond identifying gaps in definitions, influencing factors, and evaluation methodologies, our study states the limitations of existing approaches and highlights the pressing need for standardized trust assessment frameworks. Additionally, we examine the disconnect between academic research and practitioner expectations, revealing key areas where trust calibration –the alignment of users’ trust level with actual reliability or trustworthiness of a system (Turner et al., 2024) – remains a challenge in real-world SE workflows. To improve these aspects, we propose concrete steps to validate trust definitions, improve trust factors, and establish trust metrics. By systematically synthesizing these findings, this work lays the groundwork for future advancements in trust-aware LLM integration, offering concrete directions for improving the reliability, interpretability, and user confidence in these systems.
We structure our investigation around three key Research Questions (RQs) to map current understanding of trust concepts in LLM4SE. Through a systematic literature review and a survey study with SE practitioners, we examine how trust, distrust, and trustworthiness are defined in the context of LLM4SE (RQ1), identify the factors that shape trust perceptions (RQ2), and analyze existing approaches to measuring trust in LLMs (RQ3). Our findings reveal that while LLMs are increasingly adopted in SE, trust remains a loosely defined and inconsistently measured construct. Existing research neglects domain-specific definitions of trust-related concepts, often borrowing from other disciplines without adaptation to SE. In addition, trust factors, such as precision, interpretability, robustness, and workflow integration, are not consistently considered in different studies, leading to a fragmented understanding of the factors driving trust in LLM4SE. Finally, trust evaluation remains an open challenge, with limited efforts to establish standardized, multi-dimensional metrics that account for the complexity of trust in SE workflows.
To our knowledge, this is the first study to systematically examine trust in LLM4SE through a literature review and a practitioner survey. Our contributions include:
  1. (1)
    We demonstrate a current understanding of trust definitions, factors that influence trust in LLM4SE, and the metrics used to evaluate trust in LLM4SE, through an extensive review of the literature, including $18$ SE-focused papers, and analysis $70$ papers from broader trust literature.
  2. (2)
    We complement our review with a survey study of $25$ participants to map current research and gather practitioners’ perspectives on trust in LLMs in SE. We highlight gaps between literature and practical scenarios and provide a roadmap for future research.
  3. (3)
    We propose precise definitions for trust, distrust, and trustworthiness in the context of LLM4SE rooted in the existing literature and perspectives of practitioners, identify key factors influencing trust in LLM4SE and group them into model-specific and user-centric factors. Furthermore, we highlight the limitations of single-item metrics and advocate for the development of comprehensive evaluation frameworks that better capture the complexity of trust in LLM4SE.
  4. (4)
    We provide an online appendix with all of our data and results to facilitate reproducibility and encourage contributions from the community to promote trust in LLMs in SE (AnonymousRepoTrustRepo, 2024).
2. Why are Trust, Distrust, and Trustworthiness relevant for researching LLMs in Software Engineering?
Misplaced trust, whether excessive or insufficient, can lead to poor decisions and unintended consequences such as security vulnerabilities, code smells, and data leaks (Thorne, 2024; Mohsin et al., 2024; Balayn et al., 2024). For instance, researchers recently discovered that AI-powered code assistants, including GitHub Copilot and ChatGPT, often generate insecure code snippets that developers often failed to recognize as flawed (Sandoval et al., 2023; Pearce et al., 2025). The study also found that when LLM-generated suggestions contained subtle security vulnerabilities, developers were more likely to trust and use them without sufficient scrutiny, leading to increased security risks (Pearce et al., 2025). This tendency to over-rely on LLM-generated code without thorough verification goes beyond individual mistakes: it reflects a growing shift in how developers interact with LLMs and make critical coding decisions. As LLMs become integral to software development, they do more than suggest code; they influence workflows, assumptions, and even risk perception. The growing reliance on LLMs in SE requires a structured approach to understanding trust, distrust, and trustworthiness (Huang et al., 2024; Lo, 2023).
The structured understanding of trust in LLM4SE can draw inspiration from research in other fields such as HCI and AI ethics. Studies in these fields highlight key factors that influence trust, including accuracy, interpretability, and transparency (Glikson and Woolley, 2020; Sharma et al., 2024). However, trust in SE is particularly complex due to the technical nature of the domain, the potential for hidden errors in generated code, and the need for alignment with best practices (Cui et al., 2024; Balayn et al., 2024). Unlike traditional software tools, LLMs produce non-deterministic outputs, meaning the same prompt may yield different responses, making trust assessment even more challenging (Zhu et al., 2024; Ye et al., 2023). Recognizing these challenges, researchers have called for a deeper investigation into trust in SE, e.g., D. Lo (Lo, 2023) emphasizes that as software development moves toward Software 2.0, where LLMs play a central role in coding, ensuring trustworthiness in AI-assisted development is imperative.
While these calls to research highlight the urgency of understanding trust in LLM4SE, existing studies have yet to comprehensively address this issue. Despite the significance of trust in AI-powered SE tools, existing research on LLM4SE has largely overlooked definitions, factors, and metrics of trust. Our review identifies three key challenges related to trust in software engineering research. First ($CH_{1}$), only a small fraction of SE studies explicitly define trust, and distinctions between trust, distrust, and trustworthiness remain underexplored, making it difficult to assess their role in real-world SE workflows. Second ($CH_{2}$), while the broader trust literature provides valuable insight into the factors that influence trust, these perspectives have not been adequately integrated into SE research, leaving a gap in understanding what affects trust in developer interactions with LLMs. Third ($CH_{3}$), there is a noticeable absence of standardized metrics for measuring trust, and the perspectives of software practitioners on how trust should be quantified and evaluated are not well represented.
To address these challenges, we conduct a structured literature review, analyze trust research beyond SE, and complement our findings with a practitioner survey. In doing so, our goal is to provide a comprehensive understanding of trust in LLM4SE and lay the groundwork for future research and the development of trustworthy LLM-powered coding assistants aligned with developers.
3. Research Questions
This study investigates trust in LLM4SE through three RQs. Each question is addressed using a systematic review of trust literature in LLM4SE and a survey with practitioners. To provide additional context, each research question is complemented with an analysis of the underlying trust concept from the broader trust literature.
  1. RQ1
    Definition of Trust: How is trust in LLMs defined and conceptualized in the context of SE? This question aims to explore how trust, distrust, and trustworthiness are defined by analyzing existing LLM4SE literature and practitioner perspectives and address $CH_{1}$. These findings are contextualized through definitions found in broader trust literature to highlight key similarities, differences, and gaps. By analyzing all these perspectives, we provide refined, evidence-based definitions that align with academic and practitioner expectations.
  2. RQ2
    Antecedents of Trust: What are the antecedents of trust (i.e. factors that may lead to trust) in LLM in the context of SE? This question investigates the key factors that lead to the formation of trust in LLM4SE, drawing from both LLM4SE literature and practitioner experiences and address $CH_{2}$. Additional context is provided by analyzing trust factors in established trust models and related domains. The analysis provides actionable recommendations for fostering trust in LLM4SE.
  3. RQ3
    Trust Metrics: How is trust in LLMs measured in the context of SE? This question focuses on evaluating existing metrics used to assess trust in LLM4SE to address $CH_{3}$. We investigate existing trust assessment methods in LLM4SE, examine how professionals evaluate trust in real-world settings, and compare these with broader trust evaluation frameworks. By integrating these perspectives, we propose a road map for developing a comprehensive and domain-specific trust metric for LLM4SE.
4. Methodology
Our approach consists of two phases: a systematic literature review, and a survey study. Additionally, we also conducted a complimentary analysis of broader trust literature to provide further context and insights.
4.1. Phase1: Systematic Literature Review
We conducted an SLR following an approach by Kitchenman et al.  (Kitchenham, 2012). Before starting our review process, we developed RQs according to Kitchenman’s guidelines. This helped us to systematically find papers that are related to our research goals. We then performed the following steps: 1) Search for Primary Studies 2) Screening 2a) Filter by title 2b) Filter by abstract 2c) Full-text Screening 3) Snowballing 4) Manual Addition 5) Data Extraction and Analysis. Fig.1 shows a visual representation of our complete pipeline. Each step in this process was conducted independently by the two authors, who met after each step to ensure alignment and resolve any differences. The third author was consulted as needed to help settle any disputes.
[…]
[PRISMA flow diagram]A PRISMA  (Page et al., 2021) flow diagram describing the stages of our review with the number of records at each stage
Figure 1. A PRISMA  (Page et al., 2021) flow diagram that describes our review process, labeled with the number of records at each stage
4.1.1. Search for Primary Studies
We focused on the period from January 1st 2002 to November 1st 2024. We chose 2002 as our starting period as this was when the method to model language using feed-forward neural networks was published by Bengio et al.  (Bengio et al., 2000).
Databases: We selected the following databases, including ACM Digital Library, IEEE Xplore, Springer Link, Google Scholar, and DBLP. We then formulated a search string to query these databases. These databases were chosen as they were the major venues in SE, Machine Learning, HCI, and DL that encompass our research goals.
Search strings: We used various combinations of search strings to get the best results. The search string (“trust” OR “distrust” OR “trustworthiness”) AND (“SE” OR “Software Engineering” OR “Software Development” OR “Code Generation” OR “Program Repair” OR ”Automatic Program Repair” OR “Software Testing” OR “Test Generation”) AND (“LLM” OR “Large Language Model” OR “LLMs” OR “Large Language Models” OR “Deep Learning” OR “Machine Learning”) provided the most comprehensive results. Alongside ”SE” and ”Software Engineering,” we included terms related to the three SE-related activities of interest (i.e., Code Generation, Program Repair, and Software Testing) to ensure the inclusion of papers from SE and its subfields.
Search strategies: We used different mechanisms to query each database. For IEEEXplore and Springer Link, we used their advanced search feature to filter by date and venues and used their website to export the first 1,000 most relevant results as an Excel sheet. However, DBLP, Google Scholar, and ACM Digital Library do not have an export feature; thus, we wrote a simple Python program to query using a REST API and appropriate filters. After searching the databases with specified search strings and eliminating duplicates, we had 4,472 potentially relevant papers from this step. The script and the breakdown of the number of papers retrieved from each database are available in our repository  (AnonymousRepoTrustRepo, 2024).
4.1.2. Filtering by inclusion criteria
We had two iterative inclusion criteria to filter out papers from the primary search.
  • Filter by title: We screened all search outcomes by reviewing the title first. In this step, any papers with titles that contained keywords ”trust” OR ”distrust” OR ”trustworthiness” OR ”SE” OR ”Software Engineering” OR ”Software Development” OR ”Code Generation” OR ”Program Repair” OR ”Automatic Program Repair” OR ”Software Testing” OR ”Test Generation” OR ”LLM” OR ”Large Language Model” OR ”LLMs” OR ”Large Language Models” OR ”Deep Learning” OR ”Machine Learning” were considered for further screening. We kept all these keywords to ensure we did not lose any relevant papers. We were left with $3,186$ papers for further screening from this step.
  • Filter by abstract: Filtering by abstract was a manual effort by two authors of this paper. The abstract was systematically analyzed to confirm their relevance to our research goals. Papers that contained some indication of using trust or trust-related concepts were included in this step. Specifically, we searched for specific indicators in the abstracts, such as explicit mentions of trust, trustworthiness, or related concepts such as reliability, confidence, or user acceptance in the context of LLMs applied to SE tasks. Papers that merely mentioned LLMs or SE without a clear focus on trust were excluded. Additionally, abstracts that discussed trust in general AI or machine learning contexts without specific application to LLM4SE were also filtered out. From this step, we filtered out 3,163 papers, leaving us with 23 papers. The high number of excluded papers reflects the novelty of trust research specifically in the LLM4SE domain. Many papers may have touched on related topics but lacked the specific intersection of trust, LLMs, and software engineering that our study aimed to explore.
4.1.3. Full-text Screening
We conducted a thorough full-text review of the $23$ papers that passed the abstract filtering process. This step involved a detailed examination of each paper’s content to ensure its relevance to our research questions. We specifically looked for papers that addressed the intersection of trust-related concepts, SE, and LLMs. Papers were included if they:
  • Explicitly discussed trust, distrust, or trustworthiness in the context of LLMs for SE tasks.
  • Provided empirical evidence or theoretical frameworks related to trust in LLMs for SE.
  • Addressed challenges or proposed solutions for building trust in LLM-assisted SE processes.
Papers that only tangentially mentioned trust or LLMs without substantial focus on SE applications were excluded at this stage.
4.1.4. Snowballing and Manual Addition of Studies
To account for studies we might have missed with our primary search keywords, we performed snowballing on the resulting studies from our inclusion criteria. Papers in the references went through the same inclusion criteria as papers from our primary study. We first filtered them by title, and then by abstract. We added 6 papers from this step.
4.1.5. Exclusion criteria
In this step, we applied exclusion criteria to filter out studies that were irrelevant to our research questions. This involved a manual review of all papers to ensure their applicability. While trust has been widely studied in DL and HCI, we excluded papers that did not explicitly operationalize trust in SE. Our focus is on understanding trust in the context of SE, particularly regarding the integration and use of LLMs. Although trust research in DL and HCI offers valuable insights, it does not directly address SE-specific challenges such as workflow integration, developer-user interactions, and task-specific trust calibration. Therefore, we exclude these papers from our main analysis. However, to enrich our analysis, we incorporate relevant findings from broader trust research that, while not explicitly focused on SE, provide conceptual insights applicable to our study (see 4.3 for the methodology for selecting these articles). The finalized corpus contains 18 papers.
4.1.6. Data Extraction and Analysis
We iteratively went through collected papers to collect both general and specific information relevant to our RQs, following protocols suggested by  (Levett, 2022) to extract data. We create a dataset with the following data: authors, year of publication, RQs proposed, methodology, study population and characteristics (for user studies), the definition of trust, consequences of trust, the importance of trust, antecedents of trust, and metric used for trust.
Both authors worked independently to extract information and fill out the database. The authors then discussed any discrepancies with the third author. In the end, we obtained a high Inter-Rater Reliability (IRR), with a Cohen’s kappa(k) coefficient of 0.824.
Based on the extracted information, we derive our taxonomy to answer our RQs. We draw upon the framework and vocabulary from the MATCH model (Liao and Sundar, 2022) to derive our taxonomy to answer RQ2 and RQ3. The MATCH model presents factors that influence a user’s trust throughout the AI interaction process, deriving these insights from a comprehensive review of the literature, making it appropriate for our study. The MATCH model breaks down the trust-formation process into three main components:
  1. (1)
    Model Trustworthiness Attribute: These attributes are derived from the ABI framework (Dietz and Hartog, 2006) which includes three critical qualities. Ability refers to what the model can do, showcasing its competence and performance in SE-related tasks. Intention Benevolence addresses why the model was made, reflecting the intentions behind its development. Process Integrity involves how the model was trained, ensuring adherence to ethical standards and that the training data is free of biases.
  2. (2)
    Trust Affordance: Trust affordance refers to the exhibited characteristics of a system that indicate how it can be used, guiding users in evaluating the model’s trustworthiness. Trust affordance can manifest in three ways. AI-generated content is prediction, answers, and other outputs that can directly indicate the AI’s ability. Transparency regarding the model’s inner functioning can enhance users’ understanding of the model’s benevolence and process integrity. Interaction covers the ways users interact with AI, which can influence trust through the usability and design of the system.
  3. (3)
    Trust Processing: Based on users’ ability and motivation, they will partake in two different procedures of trust judgment-making. Systematic Processing involves a rigorous assessment of information to form rational trust judgments. Heuristic Processing: uses cognitive shortcuts to make quicker trust judgments, sometimes leading to incorrect judgment.
4.2. Phase2: Survey Study
With the approval of our Institutional Review Board (IRB), we conducted a survey study to investigate practitioners’ perceptions of trust-related concepts when using LLM4SE tasks. We chose a survey study over interviews or mixed methods to gather a broader and more diverse range of practitioner insights on trust in LLMs for SE tasks. User studies, particularly surveys, enable efficient data collection from larger samples while maintaining consistency for comparisons between participant groups. Unlike interviews, which are time-intensive and prone to interviewer bias, surveys provide anonymity, reducing social desirability bias, and ensuring honest responses. Additionally, the inclusion of open-ended questions in the survey allowed us to capture qualitative insights alongside quantitative data, offering a balance between depth and scalability suitable for our exploratory research goals. This approach aligns with established practices in trust research, improving the relevance and generalizability of our findings. In this study, we specifically explore the definitions of trust, its antecedents, and metrics, complementing our systematic review of the literature with insights from both experienced and novice practitioners. We employed purposive sampling  (Baltes and Ralph, 2021) to reach out to practitioners with varying levels of SE and machine learning expertise, including industrial experts, researchers, students, and software engineers. Participation was voluntary, with no incentives offered, and participants were informed about the voluntary nature of our study during solicitation.
4.2.1. Survey Structure
The survey consists of four different sections:
Section 1 examines the practitioner’s familiarity with using LLMs for various SE tasks. It qualifies practitioners for the rest of the survey, ensuring valid feedback by excluding those who have never used LLMs for the subsequent downstream tasks.
Section 2 combines blocks of questions investigating practitioners’ insights into trust-related concepts in Test Case Generation, Code Generation, and Program Repair. Blocks are presented in random order to avoid bias. Each block covers trust definitions, antecedents, and metrics. Only practitioners with experience in using LLMs for these tasks answer these questions.
Section 3 captures the general perception of the importance of trust, the desired metric for trust, and whether practitioners value the importance of these topics.
Section 4 collects demographic information about participants, such as sex, job title, and years of experience. The data is used to categorize trust-related concepts by expertise during data analysis.
4.2.2. Qualitative Metric
We derive survey metrics based on our SLR to complement it with participant’s perceptions. We capture trust definitions, antecedents, and metrics.
  • Definition: We ask practitioners with experience in using LLMs for specific tasks to define trust in code generation, test-case generation, and program repair.
  • Antecedents of Trust: We gather practitioners’ views on useful trust antecedents, as explained in  4.2.3, for particular downstream SE tasks, which allows us to investigate if trust antecedents are consistent across downstream tasks.
  • Trust Metrics: We collect information on how practitioners measure trust in specific downstream tasks through open-ended questions. This helps us compare practitioners’ trust measurement methods with those in the literature. We also ask about their code review process for both human-written and LLMs-generated code to identify similarities and differences.
4.2.3. Trust Antecedents Selection
The survey asks participants to select trust antecedents for each downstream task. We identified model-related and community-related attributes from SE, HCI, and ML literature  (Li et al., 2023b; EUw, [n. d.]; körber_2018; Widder et al., 2021a; Johnson et al., 2023; Ji et al., 2024). One author compiled all the documented antecedents, and then two authors independently categorized them, resolving disagreements through discussion. We selected the nine most popular antecedents for our survey, allowing participants to add any additional ones, which were not originally listed.
4.2.4. Data Collection
We reached out to $96$ potential participants with varying expertise and from different backgrounds, who were not involved in or aware of the purpose of this work. Out of these groups, $48$ participants answered the survey. However, we had to discard the $23$ participants due to the incomplete or poor quality of their responses leaving $25$ valid responses. The study was performed using Qualtrics  (noa, [n. d.]).
4.2.5. Survey Participants
The survey included 25 participants representing a diverse range of professional roles in software engineering and research. The majority identified as students (40%), followed by researchers (20%), engineers (16%), and professors (12%), with the remaining roles distributed among freelancers, managers, and other titles. Participants demonstrated varied levels of experience in software engineering: 44% reported more than 5 years of experience, 24% had 3–5 years, and 28% had 1–2 years, with only one respondent (4%) having less than a year of experience. In terms of education, the majority held graduate or professional degrees (60%), while 36% had bachelor’s degrees, and 4% had completed high school or equivalent.
When asked about their experience using LLMs for software engineering tasks, 36% of respondents reported 1–2 years of experience, another 36% had less than 1 year, and 16% had 3–5 years of experience. This distribution reflects a relatively early adoption stage of LLMs within software engineering, aligning with the novelty of the technology in this domain. Please refer to our online appendix  (AnonymousRepoTrustRepo, 2024) for more details.
4.2.6. Survey Validity
To solidify our survey, we conducted two pilot studies: one with $5$ participants and the other with $2$ participants who were not invited to the main survey. Based on the results of these pilot studies, we eliminated any leading questions and minor errors and enhanced our selection of trust antecedents. Our survey study is also driven by the findings from our literature survey. In addition, two of the authors carried out the data analysis individually to avoid bias and human error.
4.3. Complementary Analysis of Broader Trust literature
Our initial review of trust in LLMs in SE yielded only $18$ papers, revealing significant gaps in existing literature in LLM4SE. Thus, to supplement our systematic review and provide additional context for trust in LLM4SE, we conducted a complementary analysis of broader trust literature from HCI, DL, and automation. These fields provide well-established definitions, influencing factors, and evaluation metrics that can inform our understanding of trust in LLM4SE. Thus, we extracted trust-related concepts such as definitions, influencing factors, and evaluation metrics that could inform our findings within the SE domain.
4.3.1. Selection Process
The selection process for this complementary analysis began with revisiting the set of studies that had been excluded from our SE-focussed review due to their lack of direct alignment with LLMs, SE, and trust. Rather than discarding these studies entirely, we reassessed them to identify those that provided valuable insights into trust-related concepts, even if they were not specific to SE.
To refine our selection, we applied predefined inclusion criteria. Papers were considered relevant if they offered conceptual definitions of trust, examined the mechanisms of trust formation, proposed methodologies for measuring trust, or explored factors that influence trust formation. Following this process, we selected $70$ papers that provided meaningful contributions to understanding trust-related concepts.
4.3.2. Data Extraction and Analysis
To analyze the selected papers, we extracted key information, including definitions regarding trust-related concepts, factors contributing to trust formation, and methodologies for trust evaluation.
Two authors independently coded the extracted information with discrepancies resolved through discussion with the third author. A comparative analysis was conducted between the findings of this broader review and our systematic review focused on SE, highlighting similarities, key differences, and potential lessons that could be applied to trust in LLM-based SE systems.
We present our results in a structured format to ensure clarity. For each research question, we first summarize the insights from the SLR, followed by key takeaways from the survey study. Next, we analyze broader trust literature to contextualize the trust concepts even further. We then synthesize insights from all three sources to identify gaps, similarities, and key takeaways. Finally, we outline opportunities for future research, highlighting areas that require further exploration.
5. $RQ_{1}$: Definition of Trust
In this section, we present our results for RQ1 and discuss its implications.
Table 1. Trust concepts in LLM4SE
Aspect Trust Distrust Trustworthiness
Definition Willingness to rely on LLM outputs under uncertainty, based on positive expectations Active expectation of failure or risk in LLM outputs or behavior Intrinsic system qualities ensuring reliable, transparent, and ethical performance
Focus Positive expectations Negative expectations System characteristics
Nature User perception User perception System attribute
Example Using LLM-generated test cases based on high observed accuracy and interpretability Avoiding code from LLM due to prior output errors or security concerns A system offering explainable outputs and consistent performance aligned with SE standards
5.1. Phase1: SLR Results
In our corpus, only 27.8% ($n=5$ out of 18) of the reviewed papers explicitly defined trust, with all definitions borrowed from established concepts rather than introducing new ones. The two primary definitions of trust utilized in the corpus are:
  • “An attitude that an agent will achieve an individual’s goal in a situation characterized by uncertainty and vulnerability” (Lee and See, 2004a). This definition was cited by (Cheng et al., 2023; Wang et al., 2024c; Widder et al., 2021b; Wang et al., 2024c) (80%, $n=4$ of all the papers that defined trust).
  • ”The extent to which a user is confident in and willing to act based on, the recommendations, actions, and decisions of an artificially intelligent decision aid” (Madsen and Gregor, 2000). This definition appeared in  (Weisz et al., 2021) (20%, $n=1$).
Both definitions originate from social science and have been adapted to the Human-Machine context. These definitions emphasize that trust is rooted in the trustee’s vulnerable position, arising from uncertainty about whether the LLM will meet expectations and deliver desired outcomes. This inherent uncertainty motivates the trustee to place trust in the model (Vereschak et al., 2021). Trust is also characterized by the trustee’s positive expectation of the tool’s ability and intention to achieve its goals. Furthermore, trust is described as an attitude, how a person views and evaluates a particular situation, and is not a direct determinant of behavior. While trust can guide behaviors such as reliance on the LLM, this process is subject to individual, organizational, and cultural contexts (Liao and Sundar, 2022; Wang et al., 2024d).
Trust plays a crucial role in shaping the adoption and continued use of LLMs and software development tools (Brown et al., 2024). However, trust alone is insufficient; fostering ”appropriate trust” is essential to ensure that users rely on the model in ways that align with its actual capabilities (Kästner et al., 2021; Cheng et al., 2023; Wang et al., 2024c). Trust is an attitude stakeholders have towards the model, yet it does not necessarily indicate whether it is reliable. For instance, if a user’s trust level does not match the actual soundness of the system. In that case, they might either accept a flawed response or unjustly dismiss accurate output. Neither scenario is desirable.
One possible mediator between trust and appropriate trust is to design for model trustworthiness (Kästner et al., 2021). In the context of SE, a model is considered trustworthy if and only if it functions correctly and it is justified for stakeholders to trust the system (Kästner et al., 2021). This approach can help bridge the gap between user perception and system reliability, ultimately leading to more effective and responsive use of LLMs in SE practices.
5.2. Phase2: Survey Study Results
We analyzed 25 responses to open-ended questions on definitions of trust in LLMs for SE. One of the authors iteratively coded the survey responses, following an inductive approach to identify recurring themes. The other authors collaboratively verified and refined the final coding to ensure consistency and reliability. This analysis resulted in the following categories:
  • Functional Correctness and Performance: This was the most frequently mentioned theme across all downstream tasks. The participants emphasized the importance of LLMs generating code that meets the specified requirements, passes the test cases, and improves performance. For example, a participant noted: ”Fundamentally, I think that trust in this sense means that the code, test, or summary produced is accurate. That is if I produce a code snippet, the model hasn’t hallucinated any modules and the snippet will accomplish the task specified in the prompt…”. This theme was particularly significant for test case generation, with 75% of participants referencing it in their responses.
  • Understandability: Understandability emerged as an essential attribute, reflecting participants’ desire for outputs that are clear and easy to comprehend without extensive verification. This was especially emphasized for code generation tasks, where 26.32% of responses mentioned it. One participant remarked ”I would consider trust to mean that I would be comfortable including the LLM-generate code into a code review packet without feeling the need to verify the code in advance.”
  • Security and System Integrity: Participants expressed concern regarding the generated code’s security. They emphasized the critical role of ensuring that LLMs do not compromise the systems they generate code for. They underscored the importance of minimizing bugs and vulnerabilities. For instance, one of the participants defined trust in LLMs for code generation as ”Trust is relying on LLMs do not intentionally try to undermine systems they generate code for.” Interestingly, no participants mentioned security and system integrity as essential factors when defining trust in LLMs for test case generation.
  • Usability and Knowledge Enhancement: Participants valued the use of the generated code directly without careful evaluations and the potential of LLMs to provide additional knowledge. One of them stated, ”Trust in LLMS means I can use the code from LLMS directly in my study or work occasion.” Interestingly, the practitioner was a novice with minimal experience in SE and LLMs, reinforcing the importance of educating practitioners to foster calibrated trust in LLMs for SE applications.
  • Reliability and Consistency: This category reflects the emphasis participants placed on long-term performance and consistent behavior of LLMs in SE tasks. It was the least used theme across definitions. One participant stated, ”Trust in LLMs for Code Generation could be interpreted as the percentage of accurate code generated within an extended time frame of use.”
Our study reveals widespread use of LLMs in SE, with 99.8% of participants using them at least weekly. Despite the widespread use, only a moderate proportion trust the models (52.9% for code generation, 75% for test case generation, 54.5% for program repair). Additionally, while 84% believe that conceptualizing trust in LLM4SE is essential, 60% of participants either struggle to differentiate between trust and trustworthiness or fail to recognize the distinction altogether. This highlights the complexity of these concepts and their varying interpretations in public perception.
5.3. Complementary Analysis for Definition of Trust
Trust is a widely studied concept across disciplines such as HCI, DL, and automation. In AI research, two dominant perspectives on trust exist: (1) trust as a subjective human attitude and (2) trustworthiness as an intrinsic system property. The first perspective, common in HCI, defines trust as a user’s perception of an AI system’s reliability, shaped by factors such as interpretability, confidence, and perceived competence under conditions of vulnerability  (Vereschak et al., 2021; Jakesch et al., 2019; Vodrahalli et al., 2022; Huang et al., 2024; Ferrario and Loi, 2022; Perrig et al., 2023). In SE literature, trust is often framed similarly as an expectation about an LLM’s ability to function under uncertainty.
The second perspective, more common in DL and AI governance literature, views trustworthiness as an inherent system property, defined by attributes such as robustness, fairness, and security  (Kaur et al., 2022; Li et al., 2023b; Liu et al., 2021). In ML and AI auditing, trustworthiness is associated with fairness metrics and bias mitigation  (Lassiter and Fleischmann, 2024), while in automation, trust calibration ensures that user reliance aligns with actual system performance  (Lee and See, 2004a). Furthermore, some studies argue that trust is a multifaceted and context-dependent concept, cautioning against overly broad or narrow definitions that may limit its practical applicability  (Butler et al., 2004).
Despite these established perspectives, SE literature rarely distinguishes between trust and trustworthiness, leading to conceptual ambiguity. Our analysis also reveals that distrust is often overlooked, despite being a critical component of trust dynamics.
5.4. Discussion
Our findings reveal a significant gap between how trust is defined in the literature and how practitioners conceptualize it in real-world SE tasks. The review showed that most papers either lacked explicit definitions of trust or borrowed definitions from broader AI and social science research. These definitions describe trust as an attitude toward uncertainty and risk, focusing on the trustee’s perception rather than the system’s actual capabilities. However, our survey study demonstrates that practitioners define trust based on technical factors such as functional correctness, security, and usability, highlighting a fundamental disconnect between research and practice.
To bridge this gap, we propose the following definitions based on our systematic literature review, survey study, and insights from broader trust research. Table 1 shows overview of our definitions.
  1. (1)
    Trust Trust in LLM4SE refers to a practitioner’s willingness to rely on the recommendations or outputs of an LLM system for software engineering tasks, based on the belief that the system is capable, reliable, and aligned with their objectives under conditions of uncertainty. This trust is shaped by the system’s perceived and demonstrated performance, including correctness, security, and usability. This definition is supported by broader trust literature, where trust is conceptualized as a subjective attitude shaped by user perception  (Vereschak et al., 2021; Jakesch et al., 2019; Vodrahalli et al., 2022), as well as our SLR findings, where trust is often framed as expectation-driven. Additionally, our survey results reveal that practitioners define trust more concretely in terms of functional correctness, interpretability, and usability, emphasizing practical reliability over abstract expectations.
  2. (2)
    Distrust Distrust in LLM4SE is not merely the absence of trust but signifies a practitioner’s active expectation of risk, failure, or harm in the system’s outputs for a given SE task. This arises when the model is perceived as unreliable, unpredictable, or misaligned with ethical and functional goals, including security vulnerabilities and biased or unsafe recommendations. HCI research highlights that distrust can be driven by inconsistencies in AI behavior, misalignment with user needs, or previous negative experiences  (Degachi et al., 2024). Similarly, our survey results indicate that practitioners actively distrust LLMs when they exhibit frequent inaccuracies or security vulnerabilities, particularly in safety-critical SE tasks. The distinction between trust and distrust is thus crucial, as distrust involves an explicit expectation of failure rather than mere uncertainty.
  3. (3)
    Trustworthiness Trustworthiness in LLM4SE is the intrinsic quality of the LLM system that justifies a user’s trust, based on its ability to consistently perform tasks accurately, transparently, securely, and ethically. Unlike trust, which is user-dependent, trustworthiness is an objective property that can be evaluated through measurable factors such as robustness, fairness, security, and reliability. This aligns with DL research, where trustworthiness is defined through system properties such as fairness, robustness, and security  (Kaur et al., 2022; Li et al., 2023b; Zhang et al., 2024). In contrast to trust, which is a user-driven perception, trustworthiness represents an objective measure of AI system reliability. Our SLR found that trustworthiness in SE is rarely explicitly defined, leading to confusion between perceived and actual system reliability. Survey participants also struggled with this distinction, reinforcing the need to establish trustworthiness as a measurable property rather than an assumed user sentiment.
One key insight from our survey is that trust in LLM4SE is highly task-dependent. While 75% of respondents expressed trust in LLMs for test case generation, trust was lower for code generation (52.9%) and program repair (54.5%). This finding aligns with prior HCI research, which suggests that trust is dynamic and varies based on task complexity and perceived risk  (Jakesch et al., 2019). It also supports observations in DL and automation research, where trust is shaped by domain-specific considerations  (Lee and See, 2004a).
Additionally, our results indicate that practitioners often equate trust with reliability. Many survey participants described trust as something that builds over time through consistent correctness of outputs, rather than as an abstract attitude toward AI systems. This perspective is consistent with findings from automation and robotics, where trust is calibrated through ongoing system performance rather than being static  (Lee and See, 2004a). Similarly, in SE, trust in tools is linked to their reputation, verification, and ability to reduce the need for manual oversight  (Hou and Jansen, 2022; Rutinowski et al., 2024a). Our results suggest that trust in LLM4SE follows a similar pattern, reinforcing the need to incorporate both subjective and system-oriented factors in trust research.
Finally, security considerations emerged as a crucial component of how practitioners define trust, yet none of the reviewed SLR papers explicitly incorporated security into trust definitions. Several survey respondents associated trust with the model’s ability to produce safe, non-malicious code and prevent vulnerabilities. This gap highlights the need to integrate security into trustworthiness discussions, as emphasized in DL research on robust and privacy-preserving AI (Zhang et al., 2024; Li et al., 2023b). Additionally, institutional trust perspectives propose that trust in AI should extend beyond individual system performance to include regulatory oversight and transparency (Knowles and Richards, 2021). These insights could inform how trust in LLM4SE is conceptualized, particularly as AI governance frameworks evolve.
5.5. Opportunities for Future Work
Our findings suggest several critical avenues for future research. One key area that needs attention is the refinement and validation of trust-related concepts. The definitions that we have proposed are grounded in insights from literature review and practitioner perspectives. However, their applicability in real-world SE tasks requires further empirical assessment. A promising direction for future research could involve conducting longitudinal studies to examine how trust evolves as practitioners interact with LLM for SE tasks. Additionally, task-specific experiments focussing on different SE tasks, such as code generation, test case generation, and program repair, could help clarify whether trust in LLMs is highly task-dependent or whether users develop overarching heuristics that generalize across diverse SE contexts.
Another important area of future research is the study of distrust in the context of LLM4SE. While trust and trustworthiness have been fairly studied, distrust in LLM4SE remains an underexplored phenomenon. Our findings suggest that distrust is not merely the absence of trust but an active expectation of failure, risk, or potential harm. Future research should investigate how distrust develops and whether it leads to outright rejection of LLM-generated outputs or selective verification based on task complexity. Additionally, designing interventions to mitigate distrust—such as enhanced transparency, explicit failure explanations, and confidence estimations—could play a crucial role in improving user confidence in LLMs. Understanding the dynamics between trust, distrust, and corrective user behavior will be essential for designing AI-assisted development environments that align with practitioner expectations.
Lastly, there is a need to further explore the distinction between trust and trustworthiness. Our findings suggest that this distinction is not well understood by practitioners, as 60% of survey respondents struggled to differentiate between the two concepts. By exploring the interactions between trust, trustworthiness, and related constructs such as confidence, belief, and reliability, researchers could contribute to more precise conceptual frameworks for trust in AI. This would help bridge the gap between theoretical understandings of trust and the practical considerations that shape how users interact with LLMs in SE contexts.
Summary of $RQ_{1}$: Definition of Trust Concepts in LLM4SE Our study found that only 27.8% ($n=5$ out of 18) of reviewed papers explicitly defined trust, borrowing established definitions from social science and AI research rather than introducing new ones. Survey results from 25 participants revealed that practitioners conceptualize trust in LLMs based on functional correctness, security, understandability, and usability, whereas literature primarily frames it as an attitude toward uncertainty and risk. This discrepancy highlights a significant gap between theoretical definitions and real-world expectations. Furthermore, 60% of participants struggled to distinguish between trust and trustworthiness, indicating conceptual confusion. Trust in LLMs was found to be task-dependent, with higher trust levels for test case generation (75%) than for code generation (52.9%) and program repair (54.5%). We propose distinct definitions: trust as a practitioner’s willingness to rely on LLMs based on perceived reliability, distrust as an expectation of failure, and trustworthiness as an intrinsic system property ensuring reliability, security, and ethical alignment. Addressing these differences is crucial for fostering appropriate trust in LLMs for software engineering.
6. $RQ_{2}$: Antecedents of Trust
In this section, we present our results for RQ2 and discuss its implications.
Table 2. Trust Antecedent Overview
Construct Factor Survey Example References
Model-Specific Attributes
Ability Accuracy ”The confidence that the LLMs can identify and fix bugs in code correctly.” (Key et al., 2023; Lo, 2023; Johnson et al., 2023; Rabani et al., 2024)
Efficiency ”Make the program more efficient, without changing its functionality.” (Lo, 2023; Rabani et al., 2024)
Reliability ”My trust is based on the tool’s ability to consistently and reliably generate correct, understandable outputs.” (Johnson et al., 2023; Lo, 2023; Liu et al., 2024; Maninger et al., 2024)
Benevolence Intended Use ”The generated code fulfilled my requirements.” (Johnson et al., 2023; Wang et al., 2024c; Rabani et al., 2024)
Long-term Growth ”I don’t want it to take over my job.” (Johnson et al., 2023; Wang et al., 2024c)
Process Integrity Decision Process ”My trust relies on LLMs to not intentionally try to undermine systems.” (Russo, 2024; Lo, 2023; Maninger et al., 2024; Wang et al., 2024a)
Ethical Compliance ”The LLMs should maintain the privacy of data; it should not use the prompts for its training.” (Maninger et al., 2024; Lo, 2023; Johnson et al., 2023; Wang et al., 2024a)
User-centric Attributes
AI-generated Content Code ”The result is well-tested and properly takes care of corner cases.” (Maninger et al., 2024; Lo, 2023)
Test Case ”It can generate a rich set of tests.” (Weisz et al., 2021)
Program Repair ”Help me detect the bug and fix the bug, and I can confirm that the bug does exist.” (Noller et al., 2022)
Transparency Normative Metrics ”Trust in LLMs could be interpreted as the percentage of accurate code generated within an extended time frame of use.” (Key et al., 2023; Spiess et al., 2024; Lo, 2023; Widder et al., 2021b; Liu et al., 2024)
XAI Features ”I trust the tools if I can understand the snippet and how it functions for myself.” (Widder et al., 2021b; Johnson et al., 2023; Key et al., 2023; Weisz et al., 2021; Noller et al., 2022)
Interaction Customization ”It’s difficult to capture the requirements through a simple prompt.” (Perry et al., 2023; Johnson et al., 2023; Widder et al., 2021b)
Socialization ”Providing reliable answers which are accepted by the community like stack overflow or reliable static analysis tools.” (Cheng et al., 2023; Johnson et al., 2023)
6.1. Phase1: SLR Results
From the literature, we extracted a large collection of antecedents that impact human trust in using LLM4SE. We categorize these antecedents into two main groups: model-specific trustworthiness attributes, which are inherent to the model, and user-centric trust attributes, which are extrinsic to the model. Table 2 provides a comprehensive overview of the taxonomy we developed, constructed using the MATCH model  (Liao and Sundar, 2022) as a guiding framework. 61% (n=11 out of 18) of the studies focused on antecedents inherent to the model during the design phase; Conversely, 22% (n=4 out of 18) explored how attributes associated with interface interaction help deliver model trustworthiness to stakeholders; 17% (n=3 out of 18) explored both aspects.
6.1.1. Model-Specific Attribute
Model-Specific attributes encompass inherent characteristics and capabilities of the model independent of user perception. Understanding how model-specific attributes contribute to trust formation is essential for designing LLMs that foster appropriate trust, ensuring users neither overestimate nor underestimate the model’s reliability in SE tasks. In this section, we present results related to model-specific attributes.
  • Ability: Ability is an overarching term for LLMs’ performance or competence (Wang et al., 2024c). We identified the three most popular performance-related antecedents from our literature: accuracy, efficiency, and reliability. Accuracy is the degree to which the output matches the expected results in terms of correctness and precision (Key et al., 2023; Lo, 2023; Johnson et al., 2023). Efficiency refers to the software’s capability to use computational resources judiciously and effectively (Lo, 2023). Reliability is the ability of the software to function correctly and consistently under various conditions for an extended period (Johnson et al., 2023; Lo, 2023; Liu et al., 2024). Given the probabilistic nature of LLMs, models occasionally generate incorrect or insecure outputs. This inherent unreliability poses a key challenge to establishing trust in these models (Maninger et al., 2024). Reliability also includes the model’s ability to handle unexpected input gracefully and have appropriate fault tolerance (Liu et al., 2024).
  • Benevolence: A trustworthy model should prioritize acting in the best interest of users by aligning its goals with theirs (Wang et al., 2024c). This alignment can be categorized into two levels: short-term goal alignment and long-term goal alignment. In the short term, models must effectively address users’ immediate needs, which can vary among individuals in different situations. These needs may include offering clear advantages in usage (Johnson et al., 2023), improving productivity during the coding process (Wang et al., 2024c), and meeting specific functional expectations (Johnson et al., 2023). While achieving short-term goals is crucial for fostering initial trust, it alone is insufficient to ensure lasting trustworthiness.
    Beyond immediate outputs, models must align with the long-term goals of users. For example, developers often rely on LLMs to quickly and accurately achieve their short-term objectives. However, they are also concerned about maintaining and improving their programming skills over time, as excessive dependence on LLMs may lead to skill degradation  (Wang et al., 2024c). To address this concern, LLMs should incorporate educational content into their generated outputs, fostering learning and supporting developers in strengthening their expertise (Johnson et al., 2023). However, current models lack mechanisms that allow developers to clearly specify both their short-term and long-term goals, leading to a misalignment between the developer’s expectations and the outputs generated by the AI (Wang et al., 2024c).
  • Integrity: This attribute pertains to the training process of LLMs, including the development practice was safe and secure, under ethical guidelines (Maninger et al., 2024; Lo, 2023; Johnson et al., 2023). One study highlights the critical link between training data quality and model trustworthiness, emphasizing how issues such as biases, outdated information, or security vulnerabilities can propagate into the model’s outputs, directly impacting its integrity (Wang et al., 2024a). Regarding the training data, models are expected to reduce bias, which can be achieved by updating training data for new tasks and curating real-world datasets (Russo, 2024; Lo, 2023; Maninger et al., 2024).
6.1.2. User-Centric Attributes
User-Centric Trust Attributes are trust factors that arise from a user’s interaction with the system, including transparency, explainability, usability, and the ability to align the results with user requirements and expectations. In an ideal scenario, model-specific attributes are effectively communicated to users. However, in typical usage, users cannot directly assess a model’s trustworthiness; instead, they infer it through various affordances. Designing a trustworthy model involves creating trust affordances that deliver the actual capability of the model, so users can form appropriate trust without overestimating or underestimating the model’s capability. In this section we, we present results related to user-centric attributes.
  • AI-generated Content: AI-generated content is directly produced by LLMs. Besides natural language, this content might include SE-related outputs such as code, test cases, and program repairs, based on users’ specific needs. As AI-generated content is the primary indicator that users use for the model’s ability, the output should be well-delivered to users to receive appropriate trust assessment from them. However, our corpus has shown that the current presentation mediums for the model have difficulty capturing code structure and semantics (Maninger et al., 2024). LLM should delve into finding new mediums of representation, such as graphical representations of the generated output, to boost the effectiveness of content delivery (Lo, 2023; Weisz et al., 2021; Noller et al., 2022).
  • Transparency: Transparency is an affordance employed by developers to signal to stakeholders that their trust in the system is well-founded (Kästner et al., 2021). By offering clear justifications, transparency helps users believe that the system operates as intended. In our analysis, normative metrics and XAI features are the most emphasized factors for achieving transparency.
    One way transparency is conveyed is through normative metrics, which refer to standardized benchmarks that provide measurable, quantitative evidence of models’ performance. These metrics play a key role in evaluating the model’s ability by offering objective indicators of how well they meet expectations for accuracy, efficiency, and reliability. Importantly, these metrics help align user perceptions of the model with its actual trustworthiness  (Li et al., 2023a; Kowald et al., 2024). For example, accuracy can be assessed through metrics such as the proportion of correct results within the top N suggestions  (Lo, 2023). A high accuracy score suggests that users can confidently rely on the model to provide the correct answer within a few searches. On the other hand, Initial False Alarm (IFA) measures the number of false positives before a true positive is found. A low IFA value is preferred, as it indicates that users do not need to sift through many incorrect results  (Lo, 2023; Key et al., 2023).
    Efficiency is another key aspect evaluated through normative metrics. Efficiency metrics assess the time taken by the model to execute specific tasks, providing users with clear expectations about response times and helping to prevent surprises. This transparency about response time also builds confidence in the model’s practical usability. Similarly, reliability is often measured by the mean time between failures (MTBF), assuring users that the model consistently delivers accurate outputs without frequent errors (Widder et al., 2021b).
    Normative metrics are closely tied to transparency because they offer users the data they need to gauge the model’s performance. A well-calibrated model, for instance, provides probabilistic estimates that align closely with actual outcomes, enhancing trust by showing that the model’s predictions are accurate, effective, and reliable  (Key et al., 2023; Spiess et al., 2024). However, the current models suffer from issues like data duplication, leading to over-optimistic results  (Liu et al., 2024).
    Alongside normative metrics, XAI features serve as another essential mechanism for enhancing transparency. XAI enhances transparency by offering insights into how and why decisions are made, helping users understand the model’s reasoning and inner workings. Relevant documentation and training materials that clarify the rationale behind the model’s outputs can further aid in building trust  (Widder et al., 2021b). Studies have shown that providing clear comments and specifications in code helps users understand a program’s behavior  (Johnson et al., 2023).
    Additionally, when a model links outputs to specific source code, methodologies, and reasoning processes, users can trace the model’s logic, ensuring higher levels of provenance transparency. Finally, incorporating formal test cases and realistic test values further enhances transparency, ensuring that the model adheres to ethical guidelines and regulatory standards, which reinforces trust in the model’s governance  (Key et al., 2023; Weisz et al., 2021; Noller et al., 2022).
  • Interaction: Interaction plays a pivotal role in how users evaluate the trustworthiness of LLMs. This encompasses direct engagement with the model itself and indirect interaction through the broader user community.
    When users engage directly with the model, they often assess its customizability—the extent to which the model adapts to user preferences and feedback. Studies show that users value models that can evolve based on their input, reinforcing perceptions of the model’s ability and responsiveness  (Johnson et al., 2023; Widder et al., 2021b). Despite these affordances, limitations in prompt comprehension remain a significant barrier. Users who can effectively articulate and refine their prompts tend to achieve better outcomes, while novice users often struggle with the learning curve of prompt engineering  (Perry et al., 2023).
    Beyond direct engagement, users frequently turn to the broader community to inform their trust judgments. Socialization—interacting with user communities and leveraging collective experiences—provides valuable insights into the model’s strengths and limitations  (Cheng et al., 2023). Observing community feedback allows users to gain a more comprehensive understanding of the model’s capabilities, often supplementing the limited perspective they might form through personal use (Cheng et al., 2023). Positive examples, shared strategies, and troubleshooting tips from others help build trust by demonstrating the model’s practical value across diverse scenarios (Cheng et al., 2023).
    However, this reliance on community signals carries inherent risks. Inaccurate or misleading feedback can skew user expectations, leading to either misplaced trust or unwarranted skepticism. Ensuring that community-driven signals are evidence-based and moderated is critical to maintaining trustworthiness  (Cheng et al., 2023; Johnson et al., 2023).
6.2. Phase2: Survey Study Results
In the survey study, we asked the respondents to select one or more antecedents that would influence their trust in LLM4SE tasks. These factors can be categorized into:
  • Model attributes: Accuracy, Robustness, Ethicality, Interpretability, Controllability;
  • HCI attributes: Source Reputation, Workflow Integration, Endorsement, Community Engagement.
[…]
[Trust Antecedents]Trust Antecedents Selection Based on SE Tasks
Figure 2. Trust Antecedents Selection Based on SE Tasks
Fig. 2 is a grouped bar chart that visualizes the selection of trust antecedents across three tasks. In general, we observe that model trustworthiness attributes are valued more than human-model interaction attributes, regardless of the SE tasks. Accuracy is the most selected trust antecedent across all tasks (94% for code generation; 100% for test case generation and program repair). However, preferences for the second most important antecedents vary by task: for code generation, controllability (58.8%) and robustness (58.8%) are equally valued; for test-case generation, controllability (75%) is preferred; and for program repair, interpretability (72.7%) is valued.
In the context of program repair, one respondent added to our antecedent collection, expressing concern that their programming skills might degrade if they rely too heavily on the model. This highlights the user’s requirement for the model to support their long-term improvement in SE over solving immediate queries. It indicates the user’s emphasis on the model’s benevolence.
[…]
[Trust Antecedents]Trust Antecedents Selection Based on Years of Experience with LLMs
Figure 3. Trust Antecedents Selection Based on Years of Experience with LLMs
Fig. 3 is a heat map showing the relationship between user experiences using LLM4SE and their trust antecedent selection. The percentage is calculated as the weighted average of respondents across three SE tasks.
The visualization shows accuracy and robustness are highly valued across all experience levels. Conversely, community-associated factors (community engagement, endorsement, source reputation) are not valued by new users but gain attention with increased experience, though still less than model-related attributes. This suggests that users base trust more on direct model interaction than external validation.
A clear trend also emerges based on user experience: experienced users place less emphasis on the model’s controllability and ethical compliance. This could imply that with increasing familiarity, users developed a level of trust that mitigates concerns about how to use the tool effectively and the model’s benevolence. On the other hand, experienced users increasingly value the interpretability of the generated output and prefer models that seamlessly integrate into their established procedure.
[…]
[Trust Antecedents]Trust Antecedents based on years of Experience in SE
Figure 4. Trust Antecedents Selection Based on Years of Experience in SE
We observe a similar result from Fig. 4, a heat map between users’ year of expertise in SE and their trust antecedent selection. However, we find that there is no significant relationship between higher levels of expertise in SE and their selection of community-related factors. Interestingly, we observe that intermediates pay significantly less emphasis on robustness compared to novices and experts.
6.3. Complementary Analysis of Trust Antecedents
Trust in AI systems, including LLMs, is shaped by multiple interrelated factors, spanning technical performance, user perception, and ethical considerations. In AI research, accuracy is widely recognized as a fundamental trust factor across domains  (Khandelwal et al., 2020; Hendrycks et al., 2021; Yin et al., 2023). In SE, ensuring accuracy and reliability is essential to preventing security vulnerabilities  (Hou and Jansen, 2022; Ahuja et al., 2020; Hassan et al., 2024; Baldassarre et al., 2024a; Serban et al., 2021; Rutinowski et al., 2024b). However, trust is not solely dependent on accuracy; overconfidence in the accuracy of the model can hinder trust by making unreliable outputs appear convincing  (Huang et al., 2023). This phenomenon is observed in various AI applications, where users may develop misplaced trust in systems despite their inherent unreliability  (Banovic et al., 2023).
Beyond accuracy, trust is also influenced by the model’s ability to balance technical performance with inclusivity and multilingual capabilities. NLP research highlights that ensuring unbiased and culturally aware outputs remains a challenge when optimizing for high performance across different linguistic contexts  (Cabrero-Daniel and Sanagustín Cabrero, 2023; Zhang et al., 2023; Wang et al., 2024b). This trade-off between precision and inclusivity is particularly relevant for SE, where domain-specific terminology may not generalize well across different programming cultures.
Robustness is another critical factor in trust formation. AI models, including LLMs, often struggle with maintaining consistent performance across similar inputs. Studies show that even minor perturbations in prompts can lead to significantly different outputs  (Ye et al., 2023; Mehrotra et al., 2024), and evaluations such as PromptBench highlight LLMs’ sensitivity to word-level variations  (Zhu et al., 2024). Similar robustness challenges have been identified in NLP tasks, where adversarial manipulations can undermine model reliability  (Wang et al., 2022; Chen and Hsieh, 2023).
Transparency and accountability further shape trust in AI systems. Research suggests that users develop greater trust when they perceive AI as competent, benevolent, and transparent in its decision-making  (Baughan et al., 2023; Jakesch et al., 2019; Zerick et al., 2024; Gambo et al., 2024). However, the black-box nature of LLMs and corporate secrecy hinder transparency  (Wu et al., 2022; Liao and Wortman Vaughan, 2024; Langer et al., 2021). Techniques such as uncertainty expressions (e.g., “I’m not sure”) can improve trust by signaling limitations  (Kim et al., 2024), while XAI approaches offer interpretability at the potential cost of performance  (Wu et al., 2022). The regulatory landscape also plays a role, with frameworks like the EU AI Act emphasizing responsibility  (EUw, [n. d.]; Jobin et al., 2019; Floridi and Cowls, 2019; Nissenbaum, 1996; Stettinger et al., 2024), yet leaving questions of liability for AI errors unresolved  (Nissenbaum, 1996; Lassiter and Fleischmann, 2024). Proposed solutions, such as blockchain-based traceability, could enhance accountability in AI decision-making  (Zhang et al., 2024).
Finally, ethical considerations remain an essential, albeit sometimes secondary, factor in trust formation. While performance and usability often take priority, fairness in AI is critical to long-term trustworthiness. Studies reveal that LLMs can introduce social biases even in code generation tasks  (Liu et al., 2023), and research on models like Codex indicates a trade-off between fairness and performance  (Cui et al., 2024). Trust is also shaped by concerns about AI’s impact on professional development. Similar to educational AI, where trust is linked to fairness and correctness in recommendations  (Amoozadeh et al., 2024), SE users worry that overreliance on LLMs could degrade their programming skills. However, while not a primary focus for experienced users, ethical concerns remain an important aspect of long-term trust in AI  (Baldassarre et al., 2024c).
6.4. Discussion
Our findings reveal that trust in LLM4SE is shaped by an interplay of model-specific and user-centric attributes, highlighting the complexity of user confidence in LLMs. Accuracy emerged as a fundamental driver of trust. Our survey study also confirms that practitioners prioritize accuracy in LLM-generate outputs, mirroring observations from broader trust literature. However, overemphasis on accuracy without mechanisms for uncertainty expression can lead to misplaced trust, a well-documented concern in AI research. While accuracy is crucial, our findings underscore the necessity of equipping users with tools to critically assess reliability.
A key gap in existing research in LLM4SE is the lack of efforts to establish priority of trust antecedents. While efforts have been made to identify by model and user-specific antecedents of trust, there has been no systematic investigation into their relative importance to trust. Understanding which trust antecedents hold the highest priority for software engineers could enable more targeted improvements in LLM4SE, ensuring that development efforts are aligned with user expectations. This need for prioritization is evident in our survey study, where users consistently emphasized certain factors, such as accuracy, interpretability, controllability, and robustness over endorsement or source reputation. A lack of clarity on the relative importance of trust antecedents could lead development efforts to focus on less critical aspects while neglecting those that matter most to users.
Transparency and accountability remain central concerns for trust in LLM4SE. Our survey study indicates that experienced software engineers demand clearer explanations for model decisions, a concern echoed in broader trust literature. However, the opaque nature of LLMs, exacerbated by corporate secrecy, presents a significant barrier to transparency. Broader AI research suggests incorporating uncertainty expressions or explainable AI techniques can enhance trust. The importance of interpretability was clearly visible in our survey study, where it was deemed important across all downstream SE tasks.
Robustness is another critical dimension where expectations in software engineering diverge from broader AI applications. Our findings reveal that experienced software engineers emphasize the need for LLMs to maintain consistency across different programming tasks. This aligns with concerns in AI literature, where models frequently exhibit instability when faced with adversarial inputs or minor variations in phrasing. While robustness is a universal challenge for AI, software engineering demands particularly high levels of consistency due to the necessity of reproducibility in development workflows.
Ethical considerations, while not the primary concern for experienced software engineers, remain integral to trust formation in LLM4SE. Concerns about social biases in LLM-generated code and potential skill degradation among developers underscore the need for ethical safeguards. The ability of LLMs to handle diverse linguistic and cultural contexts is key to ensuring trust across the global community of Software Developers. While technical reliability is a prerequisite for trust, long-term confidence in AI systems requires ongoing attention to fairness, professional impact, and user empowerment. Although software engineers may not prioritize ethical considerations, its neglect could contribute to distrust.
6.5. Opportunities for Future Work
The results of this study present several avenues for future research aimed at enhancing the trustworthiness of LLMs in the context of SE tasks. An important avenue of research could be determining the relative importance of trust antecedents to prioritize trustworthy model development. Causal analysis, survey studies, controlled experiments, and A/B testing could help quantify the impact of different factors on trust. Identifying high-priority antecedents will enable targeted development of trustworthy LLMS for various SE tasks.
As trust in LLMs is highly dependent on their alignment with users’ needs, especially in a dynamic field like SE, future research should explore models that evolve with users’ skills and tasks. This includes integrating learning aids, offering skill-enhancement suggestions, and ensuring that LLMs help maintain or improve users’ programming capabilities over time. Addressing concerns about AI degrading human expertise will be crucial in promoting long-term trust.
Transparency and Interpretability emerged as significant factors for experienced users. Future work should focus on developing explainability and interpretability mechanisms that offer clear, actionable justifications for LLM-generated outputs in SE tasks. This will not only address the black-box nature of LLMs but also improve users’ understanding of how model decisions are made. Effective transparency mechanisms could help bridge the gap between technical complexity and user trust.
Another key area is community-driven trust signals, future research should investigate methods for moderating and validating community-driven feedback to ensure its reliability and usefulness. Moreover, incorporating adaptive feedback loops into LLMs could allow them to refine their performance based on real-time user input, fostering greater trust among diverse developer communities.
Lastly, tailored trust models could improve usability by adjusting trust mechanisms based on user expertise and task complexity. Novice users might benefit from simplified explanations and clear guidance, while experienced developers may seek more detailed insights into model behavior, particularly for tasks that demand high accuracy and robustness. Adapting trust signals dynamically could enhance confidence across different experience levels.
Summary of $RQ_{2}$: Antecedents of Trust Trust in LLM4SE tasks is shaped by model-specific attributes such as accuracy, robustness, and interpretability, as well as user-centric factors like transparency and usability. Accuracy emerged as the most critical factor, with 94–100% of respondents emphasizing its importance across tasks like code generation, test case generation, and program repair. Experienced users prioritize interpretability and seamless workflow integration, while novice users focus more on controllability and ethicality. Transparency mechanisms, such as XAI features and normative metrics, are essential for fostering trust but are hindered by the black-box nature of LLMs. Users also express concerns about over-reliance on these models potentially degrading their programming skills, highlighting the need for LLMs to support both immediate task completion and long-term skill development. Community-driven signals influence trust but require moderation to avoid misinformation, as direct interaction with models remains a stronger trust driver than external endorsements.
7. $RQ_{3}$: Trust Metrics
In this section, we present our results for RQ3 and discuss its implications.
7.1. Phase1: SLR Results
User trust in AI models is shaped by perceived cues and individual experiences, varying across groups (Brown et al., 2024). According to Liao et. al.  (Liao and Sundar, 2022), users may either strategically analyze LLM-generated output or rely on heuristics like intuition for trust decisions. The goal is to cultivate calibrated trust among stakeholders, avoiding both blind acceptance and unwarranted mistrust, ensuring that trust judgments are well-founded rather than based on flawed intuition.
When users lack the motivation or ability to conduct in-depth evaluations, they often form trust judgments based on surface-level cues or general impressions. Heuristics may sometimes be an intentional choice, but trust is frequently influenced by subtle, hard-to-articulate factors often linked to a “sixth sense”  (Widder et al., 2021b). Several heuristic-based trust factors have been identified in the literature.
The scope of the task significantly influences trust in LLM-generated suggestions. Developers were less likely to accept multi-line suggestions for smaller code changes, which are often more straightforward and may require less assistance. In contrast, larger code changes were associated with higher acceptance rates for multi-line suggestions (Brown et al., 2024). One hypothesis is that smaller changes provide developers with clearer ideas about their intended functionality, reducing the perceived need for AI assistance. Additionally, developers were significantly less likely to accept suggestions when editing test files. This may be because test code is often constrained to specific goals that AI suggestions may not align with, or because the suggestions might not adhere to local codebase practices, favoring global best practices instead (Brown et al., 2024).
On an individual level, a person’s experience in SE and LLMs, along with their disposition or mindset, influence trust judgment. Developers with recent experience using a programming language generated by LLMs are more likely to trust suggestions in that language  (Brown et al., 2024). Interestingly, novice programmers tend to exhibit higher trust in the security of machine-generated code than in human-generated code  (Perry et al., 2023). Developers trained to follow specific style guidelines for readability in a programming language are more likely to accept multi-line suggestions generated by LLMs  (Brown et al., 2024). Furthermore, users’ prior positive experiences with other models can increase their trust in new models, with their general tendency to trust further shaping their judgments  (Kästner et al., 2021)
Sociological aspects of the model and its building also play a role in trust judgment. People pay attention to the reputation of the company or institution that developed the model (Noller et al., 2022; Cheng et al., 2023; Johnson et al., 2023). The amount of institutional investment is also a factor that users believe indicates the model’s trustworthiness (Widder et al., 2021b). Furthermore, the population of the model’s users can lead to people’s trust, as a larger number of users offer people assurance that the model is under many levels of supervision, encouraging their belief in the model’s security provenance (Widder et al., 2021b).
Several functionality-associated factors are positively valued by users. The model is preferably easy to use (Widder et al., 2021b). For example, if the model’s features are intuitive or similar to other popular models’ usage, this allows users to operate within their comfort zone and reduces cognitive burden (Widder et al., 2021b). Also, the model’s ease of integration with the user’s existing workflow and alignment with the developer’s coding style directly assist in trust-building (Johnson et al., 2023). A polished interface design also leads to positive trust-building (Johnson et al., 2023).
During interactions, users often employ heuristics to evaluate the model. For example, a sense of control over the model’s outputs and autonomy in decision-making are highly valued. Users who can directly manipulate the code and make choices about the generated results are more likely to trust the model  (Johnson et al., 2023). Over time, as users accumulate experience and exposure to the model, they become better equipped to assess its outputs, further strengthening trust  (Widder et al., 2021b). The ability of the model to generate predictable and consistent results also plays a crucial role in trust. While users may be resilient to occasional errors, they tend to lose trust in models that produce unpredictable or surprising outputs  (Widder et al., 2021b).
7.2. Phase2: Survey Study Results
An overwhelming number of participants (96%) believe that multi-item metrics are a more effective indicator of trust than single-item metrics. This confirms our belief that trust is a multidimensional construct that must be captured more comprehensively.
Trust Metric and Measurement. For all three SE tasks, we observed that output accuracy is a highly valued aspect. Accuracy is the most frequently mentioned metric by our respondents; 38% of people who answered the question mentioned accuracy in their responses. The emphasis on accuracy highlights that users expect LLMs to deliver precise and dependable results, as reflected in the participant’s quote, ”If the generated output is flawed, there would be no point in using the model”. This expectation of output correctness aligns with the critical nature of SE, where flawed outputs can lead to significant issues.
Usability cited by 23.8% of participants, is the second most mentioned metric across tasks. It is assessed by the volume of correctly functioning code generated over time with minimal modifications needed. Users gauge trustworthiness by how well the model meets requirements without needing further changes. Usability encompasses comprehensive fulfillment of user needs, including high performance and alignment with specifications. Additionally, although LLMs are capable of generating large programs over a short period, this needs to be balanced by the ”amount of time lost trying to get a good response or wasted in debugging errors that it introduces.” With increased usability, less time needs to be spent on modifying the code, thus boosting overall productivity.
Another interesting result common across tasks is how people judge their trust in the model. This judgment is often based on their likelihood of accepting the generated output without fully understanding or carefully scrutinizing it. Respondents believe that their trust is related to their motivation to review the output. As their trust in LLMs increases, they are less likely to engage in systematic processing to gauge the model’s trustworthiness, including its ability, benevolence, and process integrity. Interestingly, these were mostly echoed by novice developers.
For Code Generation, several ability-related metrics were identified. Comprehensibility refers to the model’s ability to form logical reasoning through neural networks. Understandability pertains to the ease of understanding the code, with participants noting that their trust ”is largely tied to measures of my understanding of the generated code.” Benevolence is another metric mentioned by users, referring to the notion that ”the generated code is not maliciously trying to cause harm.”
For Program Repair, ethical compliance was mentioned as an important aspect by one of the respondents. They emphasized the importance of the generated repair following industry standards. The respondent also mentioned that their trust would be measured by the frequency of usage.
For Test Case Generation, one metric mentioned is comprehensiveness—the ability of the model to generate all possible test cases. Additionally, if the model can verify the user’s doubts, it would be deemed more trustworthy. As one respondent stated, ”I would trust the test cases more if they prove that my doubts about correct syntax and calling conventions do not apply.”
Reviewing Human versus LLM-Generated Code. Trust measurement procedures commonly include manual code reviews and running tests. Our survey revealed that 80.6% of respondents often review generated code, 16.6% rarely review it, and only 2.8% never review it. Participants often compared LLMs with human developers, suggesting that models behaving similarly to human experts are considered more trustworthy.
For code generation, 40% of participants distinguish between reviewing LLM-generated code and human-generated code. LLMs often generate code with unconventional syntax, style, and functionality, requiring more time to comprehend and integrate. One participant reported that LLMs occasionally generate ”meaningless code”, whereas ”code written by humans is usually functionally done, barring minor bugs”. However, participants acknowledge that LLM-generated code in smaller chunks can be quicker to read and process.
For test case generation, 25% of participants differentiate their review process between LLM-generated code and human-generated code, with LLM-generated tests requiring stricter scrutiny.
For program repair, 25% of participants distinguish between LLM-generated code and human-generated code reviews. Participants mentioned that the generated repair ”can often give wildly incorrect approaches (both in function and style).” Additionally, the model may produce false positive output, where the justification seems reasonable, but the change does not make sense. These issues necessitate thorough scrutiny of LLM-generated repairs. Conversely, users tend to automatically attribute more credibility to humans, assuming fewer significant flaws in their approach unless the individual has a reputation for errors.
7.3. Broader Analysis of Trust Metric
Measuring trust in LLMs for SE is a complex challenge, as it encompasses both technical and human-centric dimensions. Across AI research, trust assessment varies by domain, reflecting distinct priorities and risks. In SE and machine learning engineering, trust metrics are predominantly technical, emphasizing benchmark-driven evaluations and reproducibility  (Ahuja et al., 2020). Accuracy is often measured through adversarial testing frameworks such as FFT  (Cui et al., 2024), while TRUSTGPT assesses factual correctness by analyzing toxicity, bias, and value alignment  (Huang et al., 2023). Robustness is evaluated through adversarial success rates  (Wang et al., 2022), performance degradation from synonym substitutions  (Zhu et al., 2024), and handling of noisy inputs  (Wang et al., 2024b).
Accountability metrics in SE focus on tracking model versions, training data, and decision-making processes via immutable logs  (de Laat, 2018), while external audits further enhance reliability  (de Laat, 2018). Fairness, a growing concern in AI, is assessed using adversarial testing to detect biases in outputs  (Cui et al., 2024; Liu et al., 2023), with frameworks like TRUSTGPT standardizing toxicity and bias evaluations  (Huang et al., 2023).
In contrast, HCI research prioritizes user-centric trust metrics, relying on qualitative methods such as Likert scale assessments  (Baughan et al., 2023; Jakesch et al., 2019; Wu et al., 2023) and trust recovery studies  (Baughan et al., 2023). Transparency is evaluated through structured frameworks like POLARIS, which embeds transparency checks throughout the software development cycle  (Baldassarre et al., 2024b). Additionally, methods like LIME assess the fidelity of local explanations to overall model behavior, ensuring alignment between interpretability and performance  (Liu et al., 2021).
Despite these advancements, existing frameworks rarely integrate technical and user-centered trust metrics  (Baldassarre et al., 2024b). This gap is problematic, as trust cannot be fully established if AI systems are technically robust but socially distrusted, or user-friendly but functionally unreliable. Moreover, trust measurement frameworks often overlook the temporal dynamics of trust, which evolves over time based on user experience, failure rates, and behavioral consistency  (Lee and See, 2004b; Lewis et al., 2018). In domains such as robotics and education, trust is assessed longitudinally, capturing how user confidence fluctuates with repeated interactions. A similar approach is needed for SE, where trust in LLMs is influenced by ongoing exposure and reliability trends.
7.4. Discussion
Our study highlights the need for a comprehensive framework to measure trust in LLM4SE, as existing approaches fail to account for its multifaceted nature. While technical metrics such as accuracy, robustness, and accountability are crucial, our survey results indicate that trust also depends on usability, comprehensibility, and ethical considerations. Developers emphasized that trust is not solely determined by correctness but also by how seamlessly LLMs integrate into their workflows. The need for minimal debugging and rework reinforces the importance of trust metrics that capture practical usability alongside model performance.
Survey findings also emphasize that trust measurement should be multidimensional. A strong preference (96% of participants) for multi-item trust metrics aligns with broader research indicating that single-item trust measures oversimplify complex trust dynamics. This is particularly critical in SE, where small errors can lead to security vulnerabilities and software failures. A multidimensional approach—incorporating accuracy, usability, ethical compliance, and consistency—provides a more holistic assessment of trustworthiness.
Another major gap in current trust measurement frameworks is their failure to capture the temporal evolution of trust. Our findings indicate that trust in LLMs fluctuates over time, shaped by output quality and the extent to which developers feel they can rely on generated suggestions. This aligns with research in robotics and education, where trust is measured longitudinally through failure rates and behavioral consistency  (Lee and See, 2004b; Lewis et al., 2018). Future SE trust assessment frameworks should move beyond static evaluations and incorporate temporal metrics that track trust development over prolonged interactions with LLMs.
These insights suggest that an effective trust measurement framework for LLM4SE must integrate both technical and user-centric perspectives, consider multidimensional factors, and account for the evolving nature of trust. Addressing these gaps will enable more reliable and user-aligned assessments, ultimately fostering greater confidence in LLMs for software engineering tasks.
7.5. Opportunities for Future Work
The results of this study present several avenues for future research aimed at enhancing the metrics for trust in LLM4SE. Trust in LLMs is influenced by a variety of factors, including model accuracy, usability, ethical compliance, interpretibility, and robustness. Future research should focus on developing multi-item trust measurement frameworks that assess these dimensions in the context of SE. For instance, accuracy could be evaluated through correctness, alignment with user intentions, and contextual appropriateness. Similarly, usability could include measures such as integration ease, user control, and the effort required for debugging. Ethical compliance should be assessed through fairness evaluations and alignment with professional coding standards. A robust, multi-dimensional framework would provide a comprehensive understanding of trust and guide model improvements.
Trust in LLMs evolves over time, influenced by repeated interactions and exposure to the model. Future research should explore longitudinal studies that track how trust builds or erodes as users gain experience with the model. These studies could examine changes in trust based on output quality, model performance over time, and user feedback. Understanding the temporal dynamics of trust would help in designing models that foster long-term user confidence, making them more effective and reliable for continuous use in SE tasks.
Context-specific trust metrics are another crucial area for future work. Trust may be prioritized differently in tasks such as code generation, program repair, or test case generation, where each task involves distinct expectations and challenges. For example, code generation may emphasize accuracy, ethical considerations, and contextual relevance, while program repair may require metrics that account for error minimization, maintainability, and adherence to coding standards. Tailored trust metrics will ensure that the measurement framework aligns with the specific needs of each task and provides more meaningful insights into trust.
An important avenue for future research is trust calibration, which ensures that users’ trust levels align with the model’s actual performance. Over-reliance on faulty outputs can lead to critical errors, while under-trusting a model may result in missed opportunity and efficiency gains. Future research could investigate feedback mechanisms that help users understand why a model generated certain outputs or offer insights into the decision-making process. Transparency tools, such as model explanations and uncertainty indicators, would allow users to better align their trust with the model’s reliability. This would be especially valuable in high-stakes SE tasks, where the consequences of errors can be significant.
Summary of $RQ_{3}$: Trust Metrics We found that trust in LLM4SE is a multidimensional construct influenced by factors such as accuracy, usability, comprehensibility, and ethical compliance. A key finding from the user study revealed that 96% of the participants preferred multi-item trust metrics over single-item ones, emphasizing the need to capture trust comprehensively. User trust in LLMs for SE is influenced by both technical and human-centric factors. Accuracy and usability emerge as primary trust metrics, with accuracy being the most frequently cited criterion. Developers also highlighted the importance of comprehensibility and ethical compliance, particularly for tasks like program repair and test case generation. Sociological factors, such as the reputation of the model’s developers and its user base, further influence trust perceptions. The findings underscore the necessity of developing multi-dimensional trust measurement frameworks that account for both technical performance and user-centric dimensions to better evaluate and foster trust in LLMs.
8. Threats to Validity
We conducted our literature review following guidelines by Kithenham et al.  (Kitchenham, 2012), and our Survey Study using purposive sampling  (Baltes and Ralph, 2021). However, limitations may exist in our search strategy, data analysis, and participant recruitment.
External Threats. One potential threat to the generalizability of our results is our search strategy. Our search string might have missed some studies. To mitigate this threat, we tried multiple search strategy: 1) ”Large Language Model” OR ”Software Engineering” OR ”Trust”, 2) ”Software Engineering” OR ”SE” OR ”LLM” OR ”Large Language Model” OR ”Trust”, 3) ”Large Language Model” AND ”Software Enginnering” AND ”Trust”, 4) ”LLM” OR ”SE” OR ”Trust” OR ”Distrust” OR ”Trustworthiness”, 5) ”LLM” OR ”Large Language Model” OR ”SE” OR ”Software Engineering” OR ”Trust” OR ”Distrust” OR ”Trustworthiness”, 6) (“trust” OR “distrust” OR “trustworthiness”) AND (“SE” OR “Software Engineering”) AND (“LLM” OR “Large Language Model” OR “LLMs” OR “Large Language Models” OR “Deep Learning” OR “Machine Learning”), 7) (“trust” OR “distrust” OR “trustworthiness”) AND (“SE” OR “Software Engineering” OR “Software Development” OR “Code Generation” OR “Program Repair” OR ”Automatic Program Repair” OR “Software Testing” OR “Test Generation”) AND (“LLM” OR “Large Language Model” OR “LLMs” OR “Large Language Models” OR “Deep Learning” OR “Machine Learning”). Search string 7 returned the most results. Although this increased the time required to filter out unrelated papers, it reduced selection bias due to the large pool of results. Another threat might arise from our inclusion and exclusion criteria. To remove biases, we predefined our criteria and conducted the methodology independently by two authors. Additionally, we also conducted complementary analysis of $70$ papers that were filtered out from our exclusion criteria but contained trust concepts from broader literature.
While our user study consists of $25$ practitioners, it is important to note that this reflects a highly specialized group of individuals with direct experience in LLMs for SE tasks. Given the niche nature of LLM applications in SE, identifying qualified participants is inherently challenging. Similar SE studies have successfully used small, targeted samples (Krishna et al., 2024; coteallard2019deeplearningelectromyographichand, ). Additionally, filtering questions ensured participants were qualified, and only qualifying responses were used for analysis. We also published our responses in an online appendix (AnonymousRepoTrustRepo, 2024). While we acknowledge that larger-scale validation would be beneficial, this study provides an essential first step in mapping the trust landscape of LLMs in SE, laying the groundwork for future research.
Internal Threats. The internal threat stems from the derived taxonomy characterizing definitions, antecedents, and measurement of trust. To mitigate this, we followed a process inspired by open coding in constructive grounded theory  (Charmaz, 2006), where every attribute in our taxonomy was reviewed by four authors. Our taxonomy is also available in an online appendix  (AnonymousRepoTrustRepo, 2024).
9. Conclusion
The motivation for this study was centered on a fundamental question: Why is trust research essential in LLM4SE? To answer this, we conducted a comprehensive literature review, gathered practitioners’ feedback through a survey study, and analyzed broader trust literature. This approach allowed us to identify key gaps in how trust is defined, what factors shape it, and how it is measured. Our work provides a structured understanding of these concepts, mapping the current research landscape and highlighting their critical role in shaping LLM adoption in SE workflows.
Our findings reveal that trust in LLMs is neither absolute nor universal; it is highly task dependent and must be carefully calibrated. Developers who overtrust LLM-generated code risk introducing security vulnerabilities, code smells, and data leaks, while excessive distrust limits the potential benefits of automation and efficiency. We demonstrate that trust is influenced by multiple factors, including accuracy, robustness, interpretability, and ethical considerations. However, trust alone is not enough—trustworthiness, as an inherent property of an LLM, must be systematically evaluated through well-defined metrics rather than inferred from user perceptions alone.
By highlighting the need for structured trust assessment frameworks, our study lays the groundwork for future research on trust-aware LLM integration. As LLMs continue to shape modern software development, ensuring that trust is well-calibrated will be essential for maximizing their benefits while mitigating risks. Moving forward, researchers should focus on developing standardized trust metrics and improving transparency mechanisms to align LLM capabilities with developer expectations, ensuring their responsible and effective adoption in SE.